Set up and solve by either substitution or elimination.
Janet wishes to mix 30 pounds of coffee to sell for $6.00 a pound. To obtain the mixture, she will mmix coffee that sells for $5 a pound with coffee that sells for $7 a pound. How many pounds of each type of coffee should she use?
Having a really hard time with this one.

Answer by stanbon(75887)

Janet wishes to mix 30 pounds of coffee to sell for $6.00 a pound. To obtain the mixture, she will mmix coffee that sells for $5 a pound with coffee that sells for $7 a pound. How many pounds of each type of coffee should she use?
----------
Equation:
value + value = value
5x + 7(30-x) = 6*30
-------
5x + 7*30 - 7x = 6*30
-2x = -1*30
----
x = 15 lbs (amt. of $5.00 coffee needed)
---
30-x = 15 lbs (amt. of $7.00 coffee needed)
==============
Cheers,
Stan H.
==============